# Code Contribution and Credit in Science

### v1

@@ 5.2 Future Work
Despite their growing importance, the persistent underrecognition of software contributions suggests a need for structural interventions in how we conceptualize and reward scientific work. Building upon efforts like CRediT (Brand et al., 2015), future work should investigate potential policy changes to better align institutional incentives with the diverse spectrum of contributions that drive modern scientific progress. However, as the example of CRediT demonstrates, even well-intentioned taxonomies may reproduce existing hierarchies or create new forms of inequality if they fail to address underlying power dynamics in scientific communities. The challenge is not merely technical but social: creating recognition systems that simultaneously support innovation, ensure appropriate credit, maintain research integrity, and foster equitable participation in an increasingly computational scientific enterprise.
@@ 6 Acknowledgments
This research was supported by grants from the National Science Foundation’s (NSF) Office of Advanced Cyberinfrastructure (OAC-2211275) and the Sloan Foundation (G-2022-19347). We thank our anonymous reviewers for their helpful feedback. We especially thank Molly Blank, Shahan Ali Memnon, David Farr, and Jevin West for their valuable insights at multiple stages of this research.

### v2

@@ 5.2 Future Work
Despite their growing importance, the persistent under-recognition of software contributions suggests a need for structural interventions in how we conceptualize and reward scientific work. Building upon efforts like CRediT (Brand et al., 2015), future work should investigate potential policy changes to better align institutional incentives with the diverse spectrum of contributions that drive modern scientific progress. However, as the example of CRediT demonstrates, even well-intentioned taxonomies may reproduce existing hierarchies or create new forms of inequality if they fail to address underlying power dynamics in scientific communities. The challenge is not merely technical but social: creating recognition systems that simultaneously support innovation, ensure appropriate credit, maintain research integrity, and foster equitable participation in an increasingly computational scientific enterprise.
@@ 6 Acknowledgments
We thank our anonymous reviewers for their helpful feedback. We especially thank Molly Blank, Shahan Ali Memnon, David Farr, and Jevin West for their valuable insights at multiple stages of this research.
7 Author Contributions
Eva Maxfield Brown: Conceptualization, Data Curation, Formal Analysis, Investigation, Methodology, Software, Validation, Visualization, Writing - original draft, Writing - review & editing.
Isaac Slaughter: Formal Analysis, Visualization, Writing - review & editing.
Nicholas Weber: Conceptualization, Data Curation, Formal Analysis, Funding Acquisition, Investigation, Methodology, Project Administration, Resources, Supervision, Writing - original draft, Writing - review & editing.
8 Competing Interests
The authors have no competing interests.
9 Funding Information
This research was supported by grants from the National Science Foundation’s (NSF) Office of Advanced Cyberinfrastructure (OAC-2211275) and the Sloan Foundation (G-2022-19347).
10 Data Availability
The software and code used to train, evaluate, and apply the author-developer account matching model is available at: https://github.com/evamaxfield/sci-soft-models (https://doi.org/10.5281/zenodo.17401863). The software and code used to gather, process, and analyze the dataset of linked scientific articles and code repositories and their associated contributors and metadata is available at: https://github.com/evamaxfield/rs-graph (https://doi.org/10.5281/zenodo.17401960). The compiled rs-graph-v1 dataset, as well as data required for model training and evaluation, is available at: https://doi.org/10.7910/DVN/KPYVI1. Certain portions of the compiled rs-graph-v1 dataset, and the model training and evaluation datasets contain linked personal data (e.g., author names, developer account usernames) and are only available by request.